prove that 4^(n)+17 is divisible by 3
1
Expert's answer
2020-09-13T18:30:25-0400

Let's check for

"n = 1, f(1) = 4 + 17 = 21;"

which is divisible by 3.

  1. Let's assume that "f(n) =" "4^n + 17" is divisible by 3 for k, "k\in N:"
"f(k) = 4^k + 17 = 3 m, m\in N."

2. Let's prove that "f(k+1)" is divisible by 3:

"f(k+1) = 4^{k+1} + 17 = 4^k*4 +17=4(4^k +17) - 51" "= 4*3m - 51 = 3(4m-17)"

which is divisible by 3.

Thus "f(k+1)" is divisible by 3 if "f(k)" is divisible by 3. Hence, by the mathematical induction "f(n)=4^n + 17" is divisible by 3 for "n\in N."